# GuideGen: A Text-Guided Framework for Full-torso Anatomy and CT Volume Generation

### v2

@@ 3.2 Stage-II: Contrast-aware Autoencoder
where $w_{c}$ and $w_{r}$ are separately known as the window center and window radius. By definition, $h$ truncates the intensity range of $\mathbf{x}$ to $[w_{c}-w_{r},w_{c}+w_{r}]$ and rescales them to $[b,k+b]$. At each training step, we randomly sample the window center $w_{c}$ and window radius $w_{r}$ from the intensity range of $\mathbf{x}$ and use learnable $k$ and $b$ to map the truncated result back to the input space. Separately denoting the encoder and decoder halves of CAE as $\mathcal{E}_{I}$ and $\mathcal{D}_{I}$, the $l_{1}$ reconstruction loss on an input volume $\mathbf{v}$ can be defined as
$\mathcal{L}_{rec}=||h(\mathcal{D}_{I}(\mathcal{E}_{I}(\mathbf{v})))-h(\mathbf{v})||.$ (8)
@@
Aside from $\mathcal{L}_{rec}$, we follow the scheme of VQGAN [18] to introduce a quantization step $\mathbf{q}$ with its adversarial components, learnable codebook $\mathbf{c}$ and corresponding losses $\mathcal{L}_{vq}$ to the autoencoder training process. Specifically, CAE encodes an input volume $\mathbf{v}$ to its latent representation $\mathcal{E}_{I}(\mathbf{v})$, replaces each feature $\mathbf{z}$ with its nearest codebook entry $\mathbf{q}(\mathbf{z})=\mathbf{c}_{\mathbf{z}}$, and tries to recover the original volume $\mathbf{v}$ from the quantized latent vectors. For adversarial losses, we use a frame discriminator $\mathbf{D}_{f}$ that randomly chooses several planar slices from the CT volume and tries to distinguish between authentic slices and synthetic ones. Aside from this, we introduce a volume discriminator, $\mathbf{D}_{v}$, that does the same job from a volumetric viewpoint. These discriminators, separately focusing on anatomical details within a CT slice and the general structure of a CT volume, are not subject to contrast rescaling to avoid training instability and preserve the range of intensities for each organ. We also use perceptual loss based on a VGG-16 network [72, 59] $\mathcal{L}_{perc}$ for reducing artifacts in image reconstruction as in [8]. Overall, the loss function for CAE is
$\begin{split}\mathcal{L}_{CAE}=&\mathcal{L}_{rec}(h(\mathcal{D}_{I}(\mathcal{E}_{I}(\mathbf{v}))),h(\mathbf{v}))+\\ &\mathcal{L}_{vq}(\mathcal{E}_{I}(\mathbf{v}),\mathbf{c})+\\ &\mathcal{L}_{perc}(h(\mathcal{D}_{I}(\mathcal{E}_{I}(\mathbf{v}))),h(\mathbf{v}))+\\ &\mathcal{L}_{disc}^{frame}(\mathcal{D}_{I}(\mathcal{E}_{I}(\mathbf{v}))[i],\mathbf{v}[i],\mathbf{D}_{f})+\\ &\mathcal{L}_{disc}^{volume}(\mathcal{D}_{I}(\mathcal{E}_{I}(\mathbf{v})),\mathbf{v},\mathbf{D}_{v}),\end{split}$ (9)
where $\mathcal{L}_{vq}(\mathcal{E}_{I}(\mathbf{x}),\mathbf{c})=||\Omega(\mathcal{E}_{I}(\mathbf{x}))-\mathbf{c}_{\mathbf{z}}||_{2}^{2}+\alpha||\Omega(\mathbf{c}_{\mathbf{z}})-\mathcal{E}_{I}(\mathbf{x})||_{2}^{2}$ and $\mathcal{L}_{disc}^{\cdot}(\mathbf{x}^{\prime},\mathbf{x},\mathbf{D})=\log\mathbf{D}(\mathbf{x})+\log(1-\mathbf{D}(\mathbf{x}^{\prime}))$. $\Omega$ here is the stop-gradient operation.
3.3 Stage-III: Latent-guided Feature Generator
Since the encoded image and its semantic condition are not of the same shape, current practices usually downsample the mask volume for them to be consistent [8]. However, spatial relations between tumor and organs can be lost if the downsampling factor is too high. To promote preservation of semantic details and the textual guidance from medical prompts, we propose a Latent-guided Feature Generator (LFG) that operates entirely on the latent space to generate fine-grained image latents. As shown in Fig. 1(c), firstly, LFG encodes the torso semantics $\mathbf{m}_{A}$ and tumor semantics $\mathbf{m}_{T}$ divided from anatomies generated in Stage-I by several strided convolutional layers to downsample their spatial size. Between each downsampling convolution step, their semantic latents $\mathbf{z}^{A}$ and $\mathbf{z}^{T}$ are fused by attention layers
$\begin{split}\mathbf{z}^{A}&=\text{MLP}(\text{LayerNorm}(\text{Softmax}(\mathcal{Q}\mathcal{K}^{\top}/\sqrt{d})\mathcal{V}))\\ \end{split}$ (10)
@@ 3.3 Stage-III: Latent-guided Feature Generator
where $\mathcal{Q}=W^{q}(\mathbf{z}^{A})$, $\mathcal{K}=W^{k}(\mathbf{z}^{T})$, $\mathcal{V}=W^{v}(\mathbf{z}^{T})$ are attention matrices and $W^{q}$, $W^{k}$, $W^{v}$ are learnable mappings. Secondly, we use a diffusion model to approximate the distribution of image latents $\mathbf{z}_{0}$. In the forward process, $\mathbf{z}_{0}$ is gradually converted to Gaussian noise $\mathbf{z}_{T}\sim\mathcal{N}(0,\mathbf{I})$ through $T$ timesteps following a noise schedule $\beta^{\prime}_{1:T}$
@@
We follow a similar way as in Eq. 4 to use cross-attention layers to inject textual latents into our diffusion backbone and concatenate semantic latents $\mathbf{z}^{A}$, $\mathbf{z}^{T}$ to the diffusion variable $\mathbf{z}_{t}$ at each timestep. Finally, we construct the generative objective as in [27]:
$\mathcal{L}_{LFG}=\mathbb{E}_{t,\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})}\left[||\boldsymbol{\epsilon}-f_{\varphi}(\mathbf{z}_{t},\mathbf{z}^{A},\mathbf{z}^{T},\mathcal{E}_{T}(\mathbf{p}))||_{2}^{2}\right],$ (12)
where $f_{\varphi}$ is the 3D U-Net [11] diffusion backbone parameterized by $\varphi$ and $\mathcal{E}_{T}(\mathbf{p})$ is the textual latents encoded by the same text encoder described in Sec. 3.1. At inference time, $f_{\varphi}$ gradually recovers an image latent $\hat{\mathbf{z}}_{0}$ by reverse sampling from random noise, which is then decoded into a valid CT image by the decoder $\mathcal{D}_{I}$ from CAE.
@@ 4 Experiments
(1) Training: We compile multiple CT datasets to train GuideGen, including 12 public imaging datasets with their corresponding clinical information from The Cancer Imaging Archive (TCIA) [12] and one real-world, private dataset (RJ) on colorectal cancer. Their groundtruth label maps are retrieved using pretrained multi-organ and tumor segmentation networks [63, 34]. This collected dataset is divided into 4534 training and 1179 validation cases used for generative training. (2) Inference: For downstream usability evaluation, we additionally incorporate BTCV [46] and AMOS22 [37] datasets for multi-organ segmentation tasks and MSD-Lung Tumors (LU), MSD-Colon Cancer (CO) [3], KiTS21 [26] (KI) datasets for tumor segmentation. The rest of the experiments are conducted on the validation split of TCIA and RJ. We will elaborate on our experimental settings in Appendix 6.1.
Evaluation Metrics
The quality of our generated images are evaluated through the Learned Perceptual Image Patch Similarity (LPIPS), Fréchet Inception Distance (FID) and Fréchet Video Inception Distance (FVD) as per the general paradigm [23, 69]. We also use some these metrics to evaluate the quality of our masks due to a lack of dedicated metrics. Additionally, we report the Dice Similarity Coefficient (DSC) to evaluate image-mask consistency and the accuracy of several dedicated classifiers to measure alignment between image and certain feature described in the input prompt. Downstream Usability is evaluated through DSC and 95% Hausdorff Distance (HD95) as commonly used in segmentation tasks [14, 38]. We set HD95 to 1000 to avoid NaN values for valid comparison when there are no semantic predictions. We calculate the standard deviation for usability metrics and use a double sided $t$-test to determine the statistical significance of GuideGen’s improvement over its nearest competitor and report them in Appendix 6.2.
Implementation Details
All generative trainings are performed under an initial learning rate of $1\times 10^{-4}$ with a batch size of 1 per GPU on a multi-GPU(4) server and a patch size of $128\times 128\times 128$. The training processes use AdamW [49] optimizers with a momentum of 0.99, a weight decay of $1\times 10^{-5}$ and a linear learning rate scheduler to gradually decrease the learning rate to 0 over 500 epochs. For each diffusion model in TCSS and LFG, we use cosine noise schedule with 1000 timesteps and LFG uses DDIM denoising [60] of 200 timesteps to recover CT latents from Gaussian noise at a guidance scale of 2.5. Moreover, our CAE downsamples the original image by 4×4$\times$4 × and uses a codebook of 8192 entries.. At inference time, we train several segmentation networks using the official implementation of nnU-Net [34]. GuideGen itself is implemented using PyTorch [53] and all experiments are performed on 4 NVIDIA 4090 GPUs.
@@ 4.1 Main Results
Generation quality
Methods infer. model comput.cost Full Anatomy Tumor Only
time(s) param.(M) (GFLOPs) LPIPS FID LPIPS FID
MedGen3D [24] 67.2 48.8 1295.7 0.70 201 0.29 33.5
LDM [57] 19.9 115.2 314.9 0.67 98.6 0.30 69.1
Hu’s [29] 0.8 - 0.15 - - 0.32 64.8
GuideGen 54.1 49.1 2842.3 0.39 11.0 0.35 32.1
@@
Table 1: Comparison studies of different semantic generation methods of full-torso anatomies and tumor-only binary masks.
        Methods         input condition         inference         model         computational         generative metrics
        mask         text         time(s)         parameters(M)         cost(GFLOPs)         LPIPS         FID         FVD128
        Pinaya’s [54]         N         Y         70.2         196.1         441.0         0.465         92.0         2091
        GenerateCT[23]         N         Y         102.0         288.7         2167.9         0.521         151.7         4634
        MONAI-GM [55]         G         N         74.1         166.5         461.9         0.406         57.7         1890
        Zhuang’s [74]         G         N         45.5         81.2         2752.2         0.327         23.2         1094
        MedDDPM [16]         G         N         42.4         80.5         2748.8         0.354         35.6         1299
        MedSyn [64]         G         Y         24.7         115.1         1832.6         0.463         27.9         2267
        GuideGen         G         Y         26.1         196.8         491.4         0.242         20.5         785
        MONAI-GM [55]         R         N         74.1         166.5         461.9         0.393         54.6         1791
        Zhuang’s [74]         R         N         45.5         81.2         2752.2         0.335         39.1         1366
        MedDDPM [16]         R         N         42.4         80.5         2748.8         0.342         45.8         1619
        MedSyn [64]         R         Y         24.7         115.1         1832.6         0.396         50.1         1960
        GuideGen         R         Y         26.1         196.8         491.4         0.237         20.7         714
Table 2: Comparison studies of different image generation frameworks. The “$128$” superscript on FVD metric indicates we specifically consider CT volumes of 128 slices. A mask condition of ‘G’ or ‘R’ separately denotes semantic guidance from GuideGen-generated masks and real masks.
@@
(1) Mask quality: To validate that GuideGen’s discrete formulation is indeed superior to current methods based on continuous modeling, we choose MedGen3D [24], which is based on DDPM [27], and LatentDiffusion [57] as baseline methods and infuse them with textual information in the same way as ours. Their generation results are rounded to the nearest integer to represent semantic information. We also compare with the Hu’s deformation method described in [29] for tumor generation. The results are shown in Tab. 1. It is evident that GuideGen performs better to baselines by a large margin on anatomy generation. This is not surprising as the ambiguity problem mentioned in Sec. 3.1 significantly hinders the performance of MedGen3D and LDM when the number of semantic classes is large. This problem persists but is less prominent as the number of semantic classes drops, as shown by synthesizing only the tumor masks, where GuideGen maintains comparable performance to comparison methods. (2) Image quality: We compare GuideGen’s image generation quality with a series of generative methods from different organs, including Pinaya’s and MedDDPM [54, 16] for brain MRI generation, GenerateCT [23] and MedSyn [64] for lung CT generation, Zhuang’s [74] for abdominal CT generation and the generic medical generative framework MONAI-GM [55]. All of them are trained using the train split from TCIA and RJ datasets. At inference time, for methods that dictates a mask input, we compare their performances with GuideGen’s performance under both GuideGen-generated and real masks to ensure a fair comparison, as shown in Tab. 2. We observe that GuideGen consistently outperforms existing methods in image quality under all metrics. Particularly, we notice that methods based solely on text conditions usually perform worse than those on semantic masks. We speculate this is because the highly diverse CT textures and intensity ranges among organs are difficult to recover without voxel-wise guidance, which reflects the necessity of our TCSS. Moreover, we see little discrepancy between metrics computed from samples generated using GuideGen-generated masks and real-world masks, reflecting that our TCSS can provide comparable quality to real masks. We also evaluate the generation results qualitatively as shown in Fig. 2 by observing that existing methods do not adapt well to the high dynamic range of CT intensities, appearing either too bright or too dark in specific body regions. In contrast, GuideGen maintains a realistic CT intensity range across the full torso, thanks to our proposed CAE.
Generation-condition alignment
Methods DSC
@@
Table 4: Image-text alignment between generated CT and textual guidance. We report the accuracy of four dedicated classifiers on demographics and clinical features derived from the given medical prompts.
Figure 2: Qualitative results of different generation methods conditioned on the same textual prompts. Mask inputs to baseline models are generated with GuideGen. Two slices are shown per case. For better visualization, we use a CT intensity window of [-975,-225]HU for displaying the chest region (rows 1-2), and [-50,150]HU for the abdominal region (rows 3-6) [58]. “T” and “S” stands for frameworks that requires textual inputs or semantic inputs, respectively. More qualititative results can be found in Appendix 6.2.
      Dataset       Method       DSC
      Spleen       Kidneys       Liver       Sto.       Pan.       Adr.       Eso.       Aorta       IVC       Gall.       Duo.       Blad.       PV&SV       Avg.
      BTCV       Real       0.92       0.79       0.94       0.86       0.7       0.6       0.71       0.89       0.81       0.52       -       -       0.52       0.74
      5pt.       MONAI-GM       0.91       0.89       0.94       0.80       0.61       0.44       0.60       0.84       0.78       0.29       -       -       0.48       0.69
      Zhuang’s       0.90       0.88       0.95       0.83       0.65       0.54       0.69       0.88       0.82       0.49       -       -       0.56       0.74
      MedDDPM       0.92       0.90       0.95       0.87       0.66       0.54       0.68       0.88       0.84       0.39       -       -       0.49       0.74
      MedSyn       0.89       0.90       0.96       0.81       0.65       0.56       0.70       0.86       0.86       0.39       -       -       0.32       0.72
      GuideGen       0.94       0.92       0.96       0.91       0.74       0.59       0.73       0.91       0.87       0.49       -       -       0.65       0.79
      AMOS       Real       0.95       0.94       0.96       0.89       0.81       0.67       0.78       0.92       0.87       0.72       0.76       0.82       -       0.84
      5pt.       MONAI-GM       0.83       0.84       0.91       0.74       0.60       0.50       0.60       0.81       0.71       0.34       0.53       0.43       -       0.65
      Zhuang’s       0.85       0.87       0.90       0.66       0.63       0.46       0.61       0.82       0.73       0.26       0.55       0.62       -       0.66
      MedDDPM       0.86       0.89       0.90       0.74       0.61       0.46       0.61       0.86       0.76       0.43       0.55       0.60       -       0.69
      MedSyn       0.85       0.90       0.91       0.70       0.64       0.52       0.62       0.80       0.69       0.21       0.53       0.64       -       0.67
      GuideGen       0.93       0.92       0.95       0.86       0.75       0.55       0.73       0.88       0.81       0.65       0.67       0.73       -       0.79
@@
Table 5: Segmentation performance on multi-organ segmentation tasks using real or synthetic multi-organ masks from each generative framework.
@@
We will thoroughly evaluate GuideGen’s conditional alignment from three aspects while comparing with other baselines. (1) Image-mask alignment: Image-mask alignment can be quantitatively measured by comparing predicted results from pretrained segmentation models [63] on generated images with the actual masks given, with a higher alignment being a higher segmentation accuracy. We report the DSC of 10 major organs between predicted masks and real semantic condition input in Tab. 3. It can be seen that GuideGen behaves desirably for semantic correspondence, achieving a 2% gain on DSC compared to its nearest competitor. (2) Image-text alignment: To quantitatively evaluate the alignment between generated images and text prompt conditions, similar to [20, 23], we use a series of pretrained classifiers on certain features (age, gender, race, tumor location) to judge whether the generated images faithfully exhibits features specified in medical prompts. As shown in Tab. 4, images generated by our framework faithfully reflect the gender information and tumor location while outperforming other text-conditioned frameworks in preserving age-related and racial features. (3) Mask-prompt alignment: We measure mask-prompt alignment only qualitatively since little information except spatial relationships is depicted in generated masks. We give visualizations of mask-prompt alignment using mask synthesizers in MedGen3D (adapted by adding textual constraints to generation backbone as in Eq. 4) and GuideGen in Fig. 3 by varying tumor locations in our text prompt. It can be seen that our discrete formulation not only generates masks of accurate shape and structure, but also better preserves the spatial relationships between tumor and organs.
Usability in segmentation dataset synthesis
To assess the usability in synthesizing datasets for segmentation purposes, we use nnU-Net [34] to train segmentation models using samples generated by different frameworks as well as real-world samples from established downstream datasets. Sample synthesis is based on the anatomical masks and/or textual prompts in the validation split of our TCIA and RJ datasets.
Figure 3: Qualitative results of generated full-torso anatomical masks, with tumor masks masked in red. For each textual constraint specified, GuideGen can provide faithful anatomies while being visibly superior thanks to our discrete formulation.
@@
(1) Multi-organ segmentation: As shown in Tab. 4.1, segmentation models trained on samples generated from GuideGen can achieve comparable performance as those trained on real samples while outperforming those trained using synthetic samples generated with baseline frameworks, highlighting the generation quality of our generalized framework. (2) Tumor segmentation: Results for tumor segmentation (Tab. 6) show that our model can generate samples of lung tumors competitive with real ones while being more beneficial for other tumor segmentation tasks than baseline methods. Qualitative results can be found in Appendix 6.2.

### v3

@@ 3.2 Anatomy-aware HDR Autoencoder
where $w_{c}$ and $w_{r}$ are separately known as the window center and window radius. By definition, $h$ truncates the intensity range of $\mathbf{x}$ to $w_{c}-w_{r},w_{c}+w_{r}]$ and rescales them to $[b,k+b]$. At each training step, we randomly sample $w_{c}$ and $w_{r}$ from the intensity range of $\mathbf{x}$ and use learnable coefficients $k$ and $b$ to map the truncated result back to the input space. By denoting the encoder and decoder halves of our autoencoder as $\mathcal{E}_{I}$ and $\mathcal{D}_{I}$, the $l_{1}$ reconstruction loss on an input volume $\mathbf{v}$ can be defined as
[…]
@@
Aside from $\mathcal{L}_{\rm rec}$, to reduce noise in the reconstructed images, we incorporate adversarial components into our autoencoders. Specifically, we include a frame discriminator $\mathbf{D}_{f}$ that randomly chooses several planar slices from the CT volume and tries to distinguish between authentic slices and synthetic ones, and a volume discriminator $\mathbf{D}_{v}$, that does the same job from a volumetric viewpoint. These discriminators, separately focusing on anatomical details within CT slices and the general structure of CT volumes, are not subject to intensity transformations $h$ to avoid training instability and preserve the range of intensities for each organ. Furthermore, we use perceptual loss based on a VGG-16 network (zhang2018perceptual; simonyan2014very) $\mathcal{L}_{\rm perc}$ for reducing artifacts as in (chen2024towards). Overall, the loss function for an input CT volume $\mathbf{v}$ is
$\begin{split}\mathcal{L}_{\rm 2}=&\mathcal{L}_{\rm rec}(h(\mathcal{D}_{I}(\mathcal{E}_{I}(\mathbf{v}))),h(\mathbf{v}))+\\ &\mathcal{L}_{\rm perc}(h(\mathcal{D}_{I}(\mathcal{E}_{I}(\mathbf{v}))),h(\mathbf{v}))+\\ &\mathcal{L}_{\rm disc}(\mathcal{D}_{I}(\mathcal{E}_{I}(\mathbf{v}))[i],\mathbf{v}[i],\mathbf{D}_{f})+\\ &\mathcal{L}_{\rm disc}(\mathcal{D}_{I}(\mathcal{E}_{I}(\mathbf{v})),\mathbf{v},\mathbf{D}_{v}),\end{split}$ (9)
where $\mathcal{L}_{\rm disc}(\mathbf{x}^{\prime},\mathbf{x},\mathbf{D})=\log\mathbf{D}(\mathbf{x})+\log(1-\mathbf{D}(\mathbf{x}^{\prime}))$.
3.3 Latent-guided Feature Generator
@@
In the final stage, we use a diffusion model to approximate the distribution of image latents $\mathbf{z}_{0}$ encoded by our autoencoder based on the textual and semantic guidance from input medical prompt and TCSS. In the forward process, $\mathbf{z}_{0}$ is gradually converted to Gaussian noise $\mathbf{z}_{T}\sim\mathcal{N}(0,\mathbf{I})$ through $T$ timesteps following a noise schedule $\beta^{\prime}_{1:T}$
@@
As illustrated in Fig. 1(c), We follow a similar way as in our TCSS to use knowledge injection and cross-attention layers to inject textual information into our diffusion backbone and concatenate resampled semantic to the diffusion variable $\mathbf{z}_{t}$ at each timestep (Fig. 1(c)). The generative objective is constructed as in (ho2020denoising):
[…]
where $f_{\varphi}$ is the 3D U-Net (cciccek20163d) diffusion backbone parameterized by $\varphi$ and controlled by $\hat{\mathbf{m}}$ after being resampled to a desired output resolution as well as textual conditions $\mathbf{p}$. At inference time, $f_{\varphi}$ gradually recovers an image latent $\hat{\mathbf{z}}_{0}$ by reverse sampling from random noise, which is then decoded into a valid CT image by the decoder $\mathcal{D}_{I}$ described in Sec. 3.2. The generated sample pair $\{Resample(\hat{\mathbf{m}}),\mathcal{D}_{I}(\hat{\mathbf{z}}_{0})\}$ can then be used for constructing or augmenting segmentation dataset.
@@ 4 Experiments
Figure 2: Qualitative results of different generation methods conditioned on the same textual prompts. Mask inputs to baseline models are generated with GuideGen with tumor semantics masked in red (the first column). Two slices are shown per case. For better visualization, we use a CT intensity window of [-975,-225]HU for displaying the chest region (rows 1-2), and [-50,150]HU for the abdominal region (rows 3-6) (seeram2015computed). See our project page for more qualitative results.
@@
(1) Training: We trained GuideGen on a compiled CT dataset from 12 public TCIA sources (clark2013cancer) and a private dataset (RJ), which supplements scarce public colorectal cancer data. Text prompts were generated by a private medical LLM that converted structured records (TCIA) and reports (RJ) into free-text descriptions using the template: “The patient is {demographics group}. In this imaging, the patient’s condition is described as {clinical information}” For efficiency, ground-truth (GT) segmentation maps were pseudo-labeled using pre-trained networks (wasserthal2023totalsegmentator; isensee2021nnu). This dataset was split into 4534 training and 1179 validation cases. (2) Inference: For downstream task evaluation, we use BTCV (landman2015miccai) and AMOS22 (ji2022amos) for multi-organ segmentation as well as lung tumors (LU), colon cancer (CO) in MSD dataset (antonelli2022medical) and KiTS21 (heller2023kits21) for tumor segmentations. All other experiments used the validation split of our collected TCIA/RJ data.
Evaluation Metrics We use Perceptual Image Patch Similarity (LPIPS), Fréchet Inception Distance (FID, averaged on all slices from 3 axes) and Fréchet Video Inception Distance (FVD, axial slices as frames) to evaluate the generation quality (park2023learning; yu2024ct; lei2025lesiondiffusion). Additionally, we report the Dice Similarity Coefficient (DSC) to evaluate image-mask consistency and the accuracy of several dedicated classifiers to measure alignment between generated CT and features described in the input prompt. Downstream segmentation performance is evaluated through DSC and 95% Hausdorff Distance (HD95).
Implementation Details All generative trainings are performed under a constant learning rate of $2\times 10^{-5}$ with a batch size of 1 per GPU, a mask size of $128^{3}$ and an image size of $256^{3}$. Training processes use AdamW (loshchilov2017decoupled) optimizers with a momentum of 0.99 and a weight decay of $1\times 10^{-5}$. We implement each diffusion model with a cosine noise schedule and 1000 timesteps using PyTorch (paszke2019pytorch). GuideGen is trained and tested on 4 NVIDIA 4090 GPUs with max VRAM usages of 23.2GB and 8.9GB separately for training and inference.
@@ 4.1 Main Results
Figure 3: (a) Qualitative results of generated full-torso anatomical masks, with tumor masks masked in red. (b) Quantitative results evaluating GuideGen’s mask-prompt alignment from two dimensions including the number of tumor and tumor location.
Mask Synthesizers Parameter Full Anatomy Tumor Only
Count(M) $downarrow LPIPS\downarrow FID\downarrow$
MedGen3D(w. text) 48.8 0.70 201 0.29 33.5
LDM 115.2 0.67 98.6 0.30 69.1
Hu’s - - - 0.32 64.8
GuideGen 51.5 0.33 7.1 0.29 27.9
@@
Table 1: Comparison of different semantic generation methods of full-torso anatomies and tumor-only binary masks. $\uparrow(\downarrow)$ indicates higher(lower) values are better. We use bold and underlined metrics to indicate best and second bests.
Methods inference cond. Parameter CT generative metrics
[…]
Pinaya’s N Y 196.1 0.465 92.0 2091
GenerateCT N Y 288.7 0.521 151.7 4634
MedSyn(text-only) N Y 115.1 0.396 50.0 2012
MAISI Y N 166.5 0.406G, 0.393R 57.7G, 54.6R 1890G, 1791R
Zhuang’s Y N 81.2 0.327G, 0.335R 23.2G, 39.1R 1094G, 1366R
MedDDPM Y N 80.5 0.354G, 0.342R 35.6G, 45.8R 1299G, 1619R
MedSyn Y Y 115.1 0.304G, 0.282R 28.1G, 26.7R 1450G, 1288R
GuideGen N Y 164.1 0.248G, 0.256R 20.2G, 19.4R 791G, 745R
Table 2: Comparison studies of different image generation frameworks. Metrics suffixed with ‘G’ or ‘R’ separately denote input semantic guidance from GuideGen-generated masks and real masks. Note that GuideGen only rely on external textual inputs at inference.
@@
Generation quality (1) Mask quality: To validate that GuideGen’s ambiguity-reducing modeling is indeed superior to current methods based on continuous formulation, we choose MedGen3D (han2023medgen3d), which is based on DDPM (ho2020denoising), and LDM (rombach2021highresolution) as baselines and infuse them with textual information with cross-attention modules. Their generation results are rounded to the nearest integer to represent semantic information. We also compare with Hu’s deformation method for tumor generation (hu2023label). The results are shown in Tab. 1. It is evident that GuideGen performs better than baselines by a large margin on anatomy generation. This is not surprising as the ambiguity problem mentioned in Sec. 3.1 significantly hinders the performance of baselines when the number of semantic classes $N$ is large. This problem becomes less prominent as $N$ drops, as shown by the last two columns, where GuideGen maintains comparable performance to baseline methods. (2) Image quality: We compare GuideGen’s image generation quality with a series of generative methods, including Pinaya’s and MedDDPM (pinaya2022brain; medicalDDPM) for brain MRI generation, GenerateCT (hamamci2025generatect) and MedSyn (xu2024medsyn) for lung CT generation, Zhuang’s (zhuang2023semantic) for abdominal CT generation and the generic medical generative framework MAISI (guo2024maisi). We make minor modifications to the baseline methods, retrain them on the same datasets as GuideGen and fix the output CT volume size of all methods at $128^{3}$ to fit within our VRAM constraints. At inference, for methods that require a mask input, we use both GuideGen-generated and real masks as conditions to ensure a fair comparison in Tab. 2. We can observe that GuideGen consistently outperforms existing methods in image quality across all metrics. Particularly, we notice that methods based solely on texts usually perform worse than those with semantic masks, as the highly diverse CT textures and intensity ranges among organs are difficult to recover without voxel-wise guidance, reflecting the necessity of our TCSS. Moreover, we see little discrepancy between metrics computed from CTs given by GuideGen-generated and real-world masks, reflecting that our TCSS can provide comparable quality to real masks. Quality results in Fig. 2 also shows that GuideGen gives the most realistic and consistent CT generations.
[…]
@@
Table 4: Image-prompt alignment between generated CT and textual inputs. We report the accuracy of 4 dedicated classifiers on demographic and clinical features derived from input medical prompts.
[…]
   Cases    Spleen    Kidneys    Liver    Sto.    Pan.    Adr.    Eso.    Aorta    IVC    Gall.    Duo.    Blad.    PV&SV    Avg.
   BTCV    Real    24    0.92    0.79    0.94    0.86    0.7    0.6    0.71    0.89    0.81    0.52    -    -    0.52    0.74
   5pt.    MAISI    200    0.91    0.89    0.94    0.80    0.61    0.44    0.60    0.84    0.78    0.29    -    -    0.48    0.69
   Zhuang’s    200    0.90    0.88    0.95    0.83    0.65    0.54    0.69    0.88    0.82    0.49    -    -    0.56    0.74
   MedDDPM    200    0.92    0.90    0.95    0.87    0.66    0.54    0.68    0.88    0.84    0.39    -    -    0.49    0.74
   MedSyn    200    0.89    0.90    0.96    0.81    0.65    0.56    0.70    0.86    0.86    0.39    -    -    0.32    0.72
   GuideGen    200    0.96    0.91    0.98    0.90    0.76    0.62    0.74    0.92    0.90    0.49    -    -    0.57    0.79
   AMOS    Real    240    0.95    0.94    0.96    0.89    0.81    0.67    0.78    0.92    0.87    0.72    0.76    0.82    -    0.84
   5pt.    MAISI    200    0.83    0.84    0.91    0.74    0.60    0.50    0.60    0.81    0.71    0.34    0.53    0.43    -    0.65
   Zhuang’s    200    0.85    0.87    0.90    0.66    0.63    0.46    0.61    0.82    0.73    0.26    0.55    0.62    -    0.66
   MedDDPM    200    0.86    0.89    0.90    0.74    0.61    0.46    0.61    0.86    0.76    0.43    0.55    0.60    -    0.69
   MedSyn    200    0.85    0.90    0.91    0.70    0.64    0.52    0.62    0.80    0.69    0.21    0.53    0.64    -    0.67
   GuideGen    200    0.95    0.92    0.95    0.90    0.70    0.52    0.73    0.88    0.82    0.60    0.67    0.72    -    0.78
@@
Table 5: Segmentation performance on multi-organ segmentation tasks using real or synthetic data from each framework.
@@
Conditional alignment (1) Image-mask alignment: Image-mask alignment can be quantitatively measured by comparing mask predictions from pretrained segmentation models (wasserthal2023totalsegmentator) on generated images with the actual masks given, with a higher alignment being a higher segmentation accuracy. We report the DSC of 10 major organs between predicted masks and input semantic condition in Tab. 3. It can be seen that GuideGen behaves desirably for semantic correspondence, achieving a 6% gain on DSC compared to its nearest competitor. (2) Image-prompt alignment: To quantitatively evaluate the alignment between generated images and text prompt conditions, similar to (gu2023biomedjourney; hamamci2025generatect), we use a series of pretrained classifiers on certain features to judge whether the generated images faithfully exhibit features specified in medical prompts. As shown in Tab. 4, images generated by our framework faithfully reflect the gender information and tumor location while outperforming other text-conditioned frameworks in preserving other textual features. (3) Mask-prompt alignment: We measure mask-prompt alignment qualitatively using mask synthesizers in MedGen3D (adapted by adding textual constraints to its generation backbone) and GuideGen in Fig. 3(a) by varying the number and location of tumor sites in our text prompt. We also quantify GuideGen’s mask-prompt alignment in Fig. 3(b). It can be seen that our categorical modeling in TCSS not only generates masks of accurate shape and structure, but also preserves the spatial relationships between tumor and organs destined in the input prompts.
@@
Downstream Usability To assess the usability in synthesizing datasets for segmentation, we use nnU-Net (isensee2021nnu) to train segmentation models with samples generated by different mask-based generative frameworks. Sample synthesis is based on the masks and/or textual prompts in the validation split of our TCIA and RJ datasets. (1) Multi-organ segmentation: As shown in Tab. 4.1, GuideGen’s synthetic samples yield segmentation performance comparable to real data and superior to samples from baseline frameworks, highlighting our method’s generation quality. (2) Tumor segmentation: For tumor segmentation (Tab. 6), GuideGen-generated lung tumors yield results competitive with real data, while on other tumor segmentation tasks, GuideGen’s generated samples prove more beneficial than those from other baselines.